# Non-equilibrium dynamics of localization phase transition in the non-Hermitian Disorder-Aubry-André model

### v2

@@ IV KZS for the driven dynamics of the non-Hermitian DAA model
Figure 5: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta=0$ under OBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
@@ IV.1 General theory of the KZS
where $\varepsilon$ can represent either $\Delta$ or $\delta$, depending on the specific situation being analyzed, $\varepsilon_{0}>0$ denotes the initial distance from the critical point at $t=0$. According to the KZS, if $|\varepsilon|>R^{1/r\nu}$ with the scaling exponent $r=z+1/\nu$, the system has enough time to adapt to changes in the Hamiltonian, maintaining adiabatic conditions. Conversely, when $|\varepsilon|<R^{1/r\nu}$, the system’s internal change rate lags behind the external parameter’s rate, indicating the system entry into the impulse region Zhong1 ; Zhong2 ; Zhong3 . Below, we choose the system size $L=610$, which is sufficiently large to disregard finite-size effects in realtime simulations. We initialize the system in its ground state with a fixed $\varepsilon_{0}$ value of 1.0.
Figure 6: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta=0$ under PBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
Apparently, there are two adjustable variables $\Delta$ and $\delta$ in the non-Hermitian DAA model. Thus, both of these relevant terms should be incorporated into the full KZS form. The evolution of the localization length $\xi$ should satisfy Xuan2023 ; Liang2024
@@
Eq. (14) and Eq. (15) should be applicable for a range of $\Delta$ and $\delta$ values in proximity to the critical region of the non-Hermitian DAA model.
@@
$\xi(\Delta,R)$ $=$ $R^{-1/r_{\Delta}}f_{6}(\Delta R^{-1/r_{\Delta}\nu_{\Delta}}),$ (16)
${\rm IPR}(\Delta,R)$ $=$ $R^{s_{\delta}/r_{\Delta}\nu_{\delta}}f_{7}(\Delta R^{-1/r_{\Delta}\nu_{\Delta}}).$ (17)
This corresponds to the situation where a dynamic quench of the system is conducted along the direction of $\Delta$, with $\delta$ fixed at 0. For the case where $\delta\neq 0$, we can also fix $\delta R^{-1/r_{\Delta}\nu_{\delta}}$, and $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}$ to verify Eq. (14) and Eq. (15).
Figure 7: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.3$ under OBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
@@ IV.2 Numerical results of driven dynamics
First, we verify the scaling function Eq. (16) and Eq. (17) with the initial state deep in the localization state by linearly varying the value of the disorder potential $\Delta$. The numerical results for $\xi$ and $\rm IPR$ as functions of $\Delta$ for various driving rates $R$ under OBC are depicted in Fig. 5(a1) and Fig. 5(b1), respectively. Observation reveals that the curves corresponding to different $R$ diverge from each other near the critical region. However, when these curves are rescaled using the critical exponents of the non-Hermitian DAA model, they converge near the critical region, as illustrated in Fig. 5(a2) and Fig. 5(b2). This convergence aligns with the predictions of Eq. (16) and Eq. (17). Additionally, in Fig. 6, the dynamics of $\xi$ and $\rm IPR$ under PBC are plotted. Similarly, the rescaled curves collapse onto each other, confirming Eq. (16) and Eq. (17).
Then, we verify the KZS forms of $\xi$ and $\rm IPR$ by fixing $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.3$, which is demonstrated in Fig. 7. We numerically compute the time evolutions of $\xi$ and $\rm IPR$ for various driving rate $R$ under OBC, with results presented in Fig. 7(a1) and Fig. 7(b1), respectively. Upon rescaling the quantities according to Eq. (14) and Eq. (15), we observe that the curves corresponding to different $R$ converge near the critical region in Fig. 7(a2) and Fig. 7(b2). Similar results for $\xi$ and $\rm IPR$ under PBC are plotted in Fig. 8. After rescaling with $R$, the rescaled curves align with each other as shown in Fig. 8(a2) and Fig. 8(b2), confirming Eq. (14) and Eq. (15).
@@
Figure 8: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.3$ under PBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
@@
Figure 9: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.3$ under OBC. Curves of $\xi$ versus $\delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
We have also numerically verified the KZS for the non-Hermitian DAA model by linearly varying the value of the quasiperiodic potential $\delta$. By fixing $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.3$, we calculate the evolution of $\xi$ and $\rm IPR$ for various driving rates $R$, starting from the initial state. By rescaling these evolution curves with respect to $R$, we observe that the rescaled curves align with each other, as illustrated in Fig. 9, confirming Eq. (14) and Eq. (15). In Fig. 10, we present the evolution of $\xi$ and $\rm IPR$ under PBC. After rescaling the curves converge onto one another, in accordance with Eq. (14) and Eq. (15). These findings demonstrate the validity of the KZS in both $\Delta$ and $\delta$ directions near the critical region of the non-Hermitian DAA model.
Furthermore, we also observe that for each of the three quenching paths we considered, there is basically no difference in the results of the dynamic numerical simulations between OBC and PBC. This phenomenon can be understood from the following two aspects. Firstly, all three dynamic evolution processes involve transitions from a localized state, passing through a critical region, and ending up in another localized state. The localized states are not sensitive to changes in boundary conditions between OBC and PBC. Secondly, for each of the three paths we considered, their energy spectra exhibit similar characteristics under both OBC and PBC. For the case where $\delta=0$ and $\Delta$ is varied, the energy spectra under both OBC and PBC are basically no difference; for the case where $\delta R^{-1/r_{\Delta}\nu_{\delta}}$ is fixed and $\Delta$ is varied, the energy spectra under both OBC and PBC are real; for the case where $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}$ is fixed and $\delta$ is varied, the energy spectrum undergoes a transition from real to imaginary under both OBC and PBC.
@@
Figure 10: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.3$ under PBC. Curves of $\xi$ versus $\delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
V summary

### v3

@@ IV.1 General theory of the KZS
where $\varepsilon$ can represent either $\Delta$ or $\delta$, depending on the specific situation being analyzed, $\varepsilon_{0}>0$ denotes the initial distance from the critical point at $t=0$. According to the KZS, if $|\varepsilon|>R^{1/r\nu}$ with the scaling exponent $r=z+1/\nu$, the system has enough time to adapt to changes in the Hamiltonian, maintaining adiabatic conditions. Conversely, when $|\varepsilon|<R^{1/r\nu}$, the internal change rate of the system lags behind the rate of the external parameter, indicating the system entry into the impulse region Zhong1 ; Zhong2 ; Zhong3 .
According to Eq. (II), when $J_{L}$, $J_{R}$, $w_{j}$ and $\phi$ remain constant, adjusting either variable $\Delta$ or $\delta$ independently can induce a localized phase transition within the system. These two variables represent two independent directions related to the phase transition, and therefore both of these parameters should be incorporated into the full KZS form. The evolution of the localization length $\xi$ should satisfy Xuan2023 ; Liang2024
@@
Eq. (8) and Eq. (9) should be applicable for a range of $\Delta$ and $\delta$ values in proximity to the critical region of the non-Hermitian DAA model. Note that for the critical region of the non-Hermitian DAA model, $\nu_{\Delta}$ and $\nu_{\delta}$ are applicable simultaneously, so replacing $r_{\Delta}$ in Eq. (8) and Eq. (9) with $r_{\delta}=z_{\delta}+1/\nu_{\delta}$, both equations are also valid. In the following content, we will only discuss the situation of $r_{\Delta}$.
@@
$\xi(\Delta,R)$ $=$ $R^{-\frac{1}{r_{\Delta}}}f_{6}(\Delta R^{-\frac{1}{r_{\Delta}\nu_{\Delta}}}),$ (10)
${\rm IPR}(\Delta,R)$ $=$ $R^{-\frac{s_{\delta}}{r_{\Delta}\nu_{\delta}}}f_{7}(\Delta R^{-\frac{1}{r_{\Delta}\nu_{\Delta}}}).$ (11)
This corresponds to the situation where a dynamic quench of the system is conducted along the direction of $\Delta$, with $\delta$ fixed at 0. For the case where $\delta\neq 0$, we can also fix $\delta R^{-1/r_{\Delta}\nu_{\delta}}$ or $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}$ to verify Eq. (8) and Eq. (9).
@@ IV.2 Numerical results of driven dynamics
To verify the scaling functions Eq. (8) to Eq. (11), we numerically solve the schrödinger equation for model  (II) by using the first-order finite difference method in the time direction. The time interval $\Delta t$ is taken as 0.001. And $\varepsilon_{0}=1.0$ , which is far enough from the critical point at $\varepsilon=0$. Below, we choose the system size $L=610$, which is sufficiently large to disregard finite-size effects in realtime simulations.
@@
Figure 4: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta=0$ under OBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\Delta_{0}=1.0$.
@@
Figure 5: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta=0$ under PBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\Delta_{0}=1.0$.
@@
Figure 6: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.3$ under OBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\Delta_{0}=1.0$.
Figure 7: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.3$ under PBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\Delta_{0}=1.0$.
First, we verify the scaling function Eq. (10) and Eq. (11). Given that the state with the lowest real part of the eigenenergy often corresponds to the stable state or long-lived state of the system and holds special significance in the process of dynamical evolution, this work will primarily focus on discussing this eigenstate. We initialize the system in the state with the lowest real part of the eigenenergy at $\Delta_{0}=\varepsilon_{0}$, which is deep in the localization state. During the quenching process, the value of the disorder potential $\Delta$ is linearly changed according to Eq. (7). The numerical results for $\xi$ and $\rm IPR$ as functions of $\Delta$ for various driving rates $R$ under OBC are depicted in Fig. 4(a1) and Fig. 4(b1), respectively. Observation reveals that the curves corresponding to different $R$ diverge from each other near the critical region, indicating that the system enters the impulse region. However, when these curves are rescaled using the critical exponents of the non-Hermitian DAA model, they converge near the critical region, as illustrated in Fig. 4(a2) and Fig. 4(b2). This convergence aligns with the predictions of Eq. (10) and Eq. (11). Additionally, in Fig. 5, the dynamics of $\xi$ and $\rm IPR$ under PBC are plotted. Similarly, the rescaled curves collapse onto each other, confirming Eq. (10) and Eq. (11).
Then, we verify the KZS forms of $\xi$ and $\rm IPR$ by fixing $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.3$, which is demonstrated in Fig. 6. We numerically compute the time evolutions of $\xi$ and $\rm IPR$, initializing the system in the state with the lowest real part of the eigenenergy at $\Delta_{0}=\varepsilon_{0}$ for various driving rate $R$ under OBC. The results are presented in Fig. 6(a1) and Fig. 6(b1), respectively. Upon rescaling the quantities according to Eq. (8) and Eq. (9), we observe that the curves corresponding to different $R$ converge near the critical region in Fig. 6(a2) and Fig. 6(b2). Similar results for $\xi$ and $\rm IPR$ under PBC are plotted in Fig. 7. After rescaling with $R$, the rescaled curves align with each other as shown in Fig. 7(a2) and Fig. 7(b2), confirming Eq. (8) and Eq. (9).
We have also numerically verified the KZS for the non-Hermitian DAA model by linearly varying the value of the quasiperiodic potential $\delta$. By fixing $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.3$, we calculate the temporal evolution of $\xi$ and $\rm IPR$ for various driving rates $R$, initializing the system in the state characterized by the lowest real part of the eigenenergy at $\delta_{0}=\varepsilon_{0}$. Upon rescaling these evolution curves with respect to $R$, we observe that the rescaled curves align with each other, as illustrated in Fig. 8. This alignment confirms the validity of Eq. (8) and Eq. (9). In Fig. 9, we present the evolution of $\xi$ and $\rm IPR$ under PBC. After rescaling the curves converge onto one another, in accordance with Eq. (8) and Eq. (9).
These findings demonstrate the validity of the KZS in both $\Delta$ and $\delta$ directions near the critical region of the non-Hermitian DAA model. Assigning 0.3 to $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}$ is arbitrary. Simulation results for another randomly chosen value (0.5), which showed similar findings, are in Appendix B.
Furthermore, we also observe that for each of the three quenching paths we considered, there is basically no difference in the results of the dynamic numerical simulations between OBC and PBC. This differs from the conclusions of non-Hermitian AA model studies. In the calculations of the non-Hermitian AA model, although both the driven dynamics under PBC and OBC follow the same KZS, their manifestations exhibit certain differences. This phenomenon can be understood from the following two aspects. Firstly, all three dynamic evolution processes involve transitions from a localized state, passing through a critical region, and ending up in another localized state. The localized states are not sensitive to changes in boundary conditions between OBC and PBC. However, for the non-Hermitian AA model, the phase transition under PBC corresponds to a local-to-extended phase transition, whereas under OBC it manifests as a local-to-skin effect phase transition. Secondly, for each of the three paths we considered, their energy spectra exhibit similar characteristics under both OBC and PBC. For the case where $\delta=0$ and $\Delta$ is varied, the energy spectra under both OBC and PBC are basically no difference; for the case where $\delta R^{-1/r_{\Delta}\nu_{\delta}}$ is fixed and $\Delta$ is varied, the energy spectra under both OBC and PBC are real; for the case where $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}$ is fixed and $\delta$ is varied, the energy spectrum undergoes a transition from real to imaginary under both OBC and PBC. For the non-Hermitian AA model, under PBC, the energy spectrum undergoes a real-to-complex transition, while under OBC, such behavior is absent.
Figure 8: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.3$ under OBC. Curves of $\xi$ versus $\delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\delta_{0}=1.0$.
Figure 9: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.3$ under PBC. Curves of $\xi$ versus $\delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\delta_{0}=1.0$.
V HKZS in the overlapping critical region of non-Hermitian DAA and Anderson localization
As detailed in Sec. IV, these three quenching paths under consideration do not traverse the skin-effect region. As previously mentioned, for $\delta<0$, an overlapping critical region emerges where the non-Hermitian DAA and non-Hermitian Anderson localization coexist. Under PBC, the energy spectra of non-Hermitian DAA systems in the $\delta<0$ region all exhibit imaginary components, making it impossible to analyze driven dynamics. However, under OBC, since the energy spectra remain real within a certain range, this provides an opportunity to investigate the driven dynamics in the overlapping region.
V.1 General theory of the HKZS
In the overlapping critical region where $\delta<0$, scaling functions of non-Hermitian Anderson transitions and the non-Hermitian DAA model are both crucial. To study scaling behavior here, a HKZS is proposed. In a typical case where the overlapping region is assumed to consist of region A and B, it hypothesizes that the driven dynamics in the overlapping region can be described by KZS of both regions simultaneously and a constraint exists between their scaling functions Xuan2022 ; Xuan2023 ; Yin2017 ; zhai2018 .
Here, we take the critical properties of $\rm IPR$ to illustrate the HKZS. In the critical region of non-Hermitian Anderson transition, the evolution of the $\rm IPR$ should satisfy
${\rm IPR}(\Delta,R)=R^{\frac{s_{A}}{r_{A}\nu_{A}}}f_{8}(\Delta R^{-\frac{1}{r_{A}\nu_{A}}}).$ (12)
where $(\nu_{A},s_{A},z_{A})=(1.99,1.99,2)$ are the critical exponents of the non-Hermitian Anderson transition Sun2024 and $r_{A}=z_{A}+1/\nu_{A}$.
According to the hybrid scaling law, both the scaling functions of $\rm IPR$, i.e., Eqs. (9) and (12), are applicable in the critical region where $\delta<0$. Combining Eqs. (9) and (12), the constraint between these two scaling functions should satisfy
$f_{5}(A,B)=A^{\kappa}f_{9}\left[B(A)^{\chi}\right]$ (13)
where $A\equiv{\delta}R^{-1/r_{\Delta}\nu_{\delta}}$, $B\equiv{\Delta}R^{-1/r_{\Delta}\nu_{\Delta}}$, $\kappa\equiv{r_{\Delta}\nu_{\delta}(s_{\delta}/r_{\Delta}\nu_{\delta}-s_{A}/r_{A}\nu_{A})}$ and $\chi\equiv{r_{\Delta}}\nu_{\delta}(1/r_{A}\nu_{A}-1/r_{\Delta}\nu_{\Delta})$. We find that $\chi$ and $\kappa$ include both the critical exponents of non-Hermitian Anderson model and non-Hermitian DAA model, which give the constraint between these scaling functions.
V.2 Numerical results of driven dynamics
Here, we numerically verify these scaling tfunctions in the overlapping critical region with $\delta<0$.
By fixing $\delta R^{-1/r_{\Delta}\nu_{\delta}}=-0.3$, Eq. (9) is firstly verified. We numerically compute the time evolutions of $\rm IPR$, initializing the system in the state with the lowest real part of the eigenenergy at $\Delta_{0}=\varepsilon_{0}$ for various driving rate $R$ under OBC. The dynamic evolution process involve transitions from a localized state, passing through a skin effect region.
Figure 10: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta R^{-1/r_{\Delta}\nu_{\delta}}=-0.3$ under OBC. Curves of $\rm IPR$ versus $\Delta$ before (a) and after (b) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\Delta_{0}=1.0$.
Figure 11: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta=-0.1$ under OBC. Curves of $\rm IPR$ versus $\Delta$ before (a) and after (b) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\Delta_{0}=1.0$.
Figure 12: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta=-0.1$ under OBC. (a) Curves of $f_{5}$ versus $\Delta$ for different $R$. (b) Curves of $A^{-\kappa}f_{5}$ versus $B(A)^{\chi}$ for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used. $\Delta_{0}=1.0$.
In Fig. 10, the scaling properties of $\rm IPR$ versus $\Delta$ for $\delta R^{-1/r_{\Delta}\nu_{\delta}}=-0.3$ are plotted. After rescaling according to Eq. (9), the rescaled curves collapse onto each other well, confirming Eq. (9). This quenching path crossing the skin-effect region reveals that the skin effect creates depression zones in the IPR curves, which suggests the emergence of a skin-effect state Zhaifphy . It exhibits certain differences from the evolutionary results of the three previously discussed quenching paths that transition from localized states to localized states. Therefore, the choice of boundary conditions does not significantly affect the applicability of the KZS, however, during the evolution process of the system, its evolutionary trend will exhibit some differences depending on whether it passes through the skin-effect region.
Next, we take a fixed $\delta<0$ and vary $R$ to verify Eq. (12). The numerical results for $\delta=-0.1$ are plotted in Fig. 11. The collapse of the rescaled curves shown in Fig. 11 (b) also confirms Eq. (12). Therefore, numerical results in Fig. 10 as well as Fig. 11 confirm the first hypothesis of the hybrid scaling law.
Finally, we also verify Eq. (13) using a fixed $\delta<0$ with varying $R$. The numerical results of $f_{5}=\rm IPRR^{-s_{\delta}/r_{\Delta}\nu_{\delta}}$ as a function of $\Delta$ for various $R$ are plotted in Fig. 12 (a). By rescaling $\Delta$ as $B(A)^{\chi}$, we find that the rescaled curves collapse very well, verifying Eq. (13) and the second hypothesis of the hybrid scaling law.
VI Experimental scheme
Recently, the validity of the KZS in non-Hermitian systems has been experimentally verified through a time-multiplexed photonic quantum walk system  xuekzs2021 . Building on this progress, our research results are also expected to be verified in similar platforms. This approach employs time-multiplexed encoding, where photon polarization and arrival time encode spatial and internal degrees of freedom. Non-Hermiticity is introduced via polarization-dependent loss, dynamically controlled by an electro-optic modulator (EOM). Our DAA model is constructed by incorporating quasi-periodic potential through position-dependent phase operators xueprl2022 and by introducing disorder via EOM-driven random modulation of coin operation angles xueNC2022 . The quantum walk dynamics are governed by the Floquet operator xueprl2022 ; xueNC2022 . The photon state is initialized and evolved through the quantum walk, with system parameters gradually varied to traverse the phase transition point. Photon distributions are recorded using avalanche photodiodes (APDs) to measure dynamic observables. Defect density and fluctuations are analyzed to validate KZS behavior, with critical exponents $\nu$ and $z$ extracted from power-law fits to confirm the universality class of the phase transition. This framework provides a platform for investigating the interplay of disorder and quasi-periodicity in non-equilibrium quantum systems, deepening insights into localization phenomena.
VII summary